please solve and check rational Equation
1) 4/x+2-1/x= 1/x

2) 15/x-4=6/x+3
3) Find the two consecutive positive even Integers whose product is 168.

Answer by lazi(15)

1) 4/x + 2 - 1/x = 1/x
3/x + 2 = 1/x
- 1/x - 2 = -1/x - 2
----------------------------
2/x = - 2
times x and divide by - 2 on both side
- 1 = x
2) 15/x - 4 = 6/x + 3
- 6/x + 4 = - 6/x + 4
----------------------------
9/x = 7
times x and divide by 7 on both side
9/7 = x
3)----168
-----/---\
----2----84
--------/--\
-------2---42
----------/--\
---------2---21
------------/ \
-----------3----7
2 * 2 * 2 * 3 * 7 = 168
2 * 2 * 3 = 12 ; 2 * 7 = 14
formula x(x-2) = 168
x^2 - 2x - 168 = 0
(x-14)(x+12) = 0
x = 14 and 12